NEXT
Make New Entry,
Make Followup Entry
User name nmuangma
Log entry time 08:10:29 on April14,2011
Entry number 355419
keyword=owl shift summary
April 13, 2011
SL: Navaphon Muangma
TO: John Leckey
Owl Shift Summary:
++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++++
Experiment: E08009
beam energy = 4.450 GeV
left HRS P0 = 3.6014 GeV/c
left HRS angle = 20.5 degs
right HRS P0 = 1.5100 GeV/c
right angle = -47.0 degs
0:00 Current scan (20,30,40, & 50) uA (3486-90)
0:38 Hall C request us to turn off the beam for their test.
1:00 Start BigBite Optic Scan run(3493-5) with magnet off and on.
1:38 resume production at 50 uA (60 uA gave unstable beam)
5:03 All HV for BigBite is OFF.
Run Summary:
Single rates runs fro current scan:
He4 target, RHRS-theta = 47
3486 @ 20 uA
3487 @ 30 uA
3488 @ 40 uA
3490 @ 50 uA
Run Summary:
Run | Events | Ibeam [uA] | BCM U1 [C] | Accumulated Q [C] | Deadtime [%] | Target | Comments |
3493 | 257 K | 4.0 uA | N/A | N/A | 5 | Carbon Optic | Optic |
3494 | 700 K | 4.0 uA | N/A | N/A | 7 | BeO | Optic: with BB magnet OFF |
3495 | 500 K | 4.0 uA | N/A | N/A | 7.5 | BeO | Optic: with BB magnet ON |
Run | Events | Ibeam [uA] | P_miss intetral | BCM U1 [C] | Accumulated Q [C] | Deadtime [%] | Target | Comments |
3496-7 | Junk | 50-60 | N/A | N/A | N/A | | | Junk |
3498 | 1.0 M | 50 uA | 29757 | 0.12890 | 0.12890 | 9 | He4-20cm | Production |
3499 | 1.0 M | 50 uA | 29885 | 0.12960 | 0.25850 | 9 | He4-20cm | Production |
3500 | 1.0 M | 50 uA | 30242 | 0.12860 | 0.38710 | 9 | He4-20cm | Production |
3501 | 1.0 M | 50 uA | 29882 | 0.12880 | 0.51590 | 9 | He4-20cm | Production |
3502 | 1.0 M | 50 uA | 29453 | 0.12880 | 0.64470 | 9 | He4-20cm | Production |
3503 | 1.0 M | 50 uA | 29550 | 0.12860 | 0.77330 | 9 | He4-20cm | Production |
3504 | 1.0 M | 50 uA | 29648 | 0.12850 | 0.90180 | 9 | He4-20cm | Production |
3505 | 1.0 M | 50 uA | | | | 9 | He4-20cm | Production |
3506 | 1.0 M | 50 uA | | | | 9 | He4-20cm | Production |
Sum BCM U1 (C) = 0.90180